2.5 g of an organic compound was dissolved in 78 g of benzene.the vapour pressure of solute and solvent were found to be 610 mm634mm of mercury respectively.calculate the molecular weight of organic compound
 

Grade:12

1 Answers

ROSHAN MUJEEB
3 years ago

HELLO SHYAM,
MOLES of benzene = 78/78
= 1
Moles of solute = 2.5/M
=>Mole fraction of benzene
= X = 1
-------
1 + 2.5/M
=> 610 = 635X
=> = 635
--------
1 + 2.5/M
=> 2.5/M = 635
------ – 1
610
=> M = 61
THANK YOU